OpenStudy (anonymous):

39. write an equation in slope-intercept formof the line through point p(6, 7) with slope -2. a)y-7=-2(x-6) b)y=-2x+7 c)y=-2x+19 d)y-6=-2(x-7)

OpenStudy (anonymous):

y=mx+b plug in the point and slope 7=(-2)6+b 7=-12+b add 12 on both sides 19=b now that you know b plug in the slope and b y=mx+b soooo.. y=-2x+19

OpenStudy (anonymous):

thanks